if 1/40,000 individuals are affected by an autosomal recessive disorder in a population in HWE, what is the probability a carrier and someone with no family history will have an affected child?
Answer #1

Answer:

q^2=1/40,000 =2.5 x 10^-5

q=0.0000158

p+q=1

p=1-q=1-0.0000158=0.9999

Heterozygous genotype is a carrier= 2pq= 2 x 0.0000158 x 0.9999 =3.15 x 10^5

Probability of carrier couple having an affected baby in autosomal recessive disorder is ¼

One of the parent is confirmed carrier, other is 3.15 x 10^5 percent chance of being a carrier

Hence the total probability is = ¼ x 3.15 x 10^5= 0.865 x 10^5